# Diffusion in driven periodic systems: from weak to strong damping regimes

### v1

@@ Abstract
We consider diffusion of an inertial Brownian particle moving in a symmetric spatially periodic potential and driven by an unbiased time periodic force of a moderate frequency. The study is focused on the dependence of its diffusion coefficient $D$ on the damping $\gamma$ ranging from the weak to strong one. We reveal the mechanisms of counterintutive effects in which diffusion can increase despite the grow of friction and exhibits both minima and maxima. Their origin is explained by analysis of the deterministic dynamics perturbed by thermal fluctuations. In the weak damping regime the maximum of diffusion is induced by a non-monotonic change of population of the running orbits when the friction increases. In the strong dissipation regime the effect is caused by bifurcation of periodic orbits oscillating symmetrically around the potential maximum and their displacement towards the potential minimum when the damping coefficient increases. It induces the change of the effective barrier of the periodic potential and its vanishing at the maximum of $D$.
I Introduction
@@
$D=\frac{k_{B}T}{\gamma},$ (1)
where $k_{B}$ is the Boltzmann constant, $T$ is temperature of the medium (thermostat) in which the particle moves and $\gamma$ is the friction (damping) coefficient which characterizes the interaction strength of the Brownian particle with the environment. The same relation is also fulfilled for a Brownian particle subjected to a constant force $F={const}$. From this equation it follows that $D$ depends inversely on the friction coefficient $\gamma$. In other words, $D$ decreases when $\gamma$ increases. Studies on e.g. surface diffusion reveals that in some regimes the more general relation is satisfied, namely,
@@
The exponent $\sigma$ is not universal but depends on a system, its model and the corresponding assumptions. Different values have been reported like $\sigma=1/2$ sigma05 , $\sigma=1/3$ sigma03 , $\sigma\in(0.64,0.76)$ sigma064 , $\sigma\in(0,1/3)$ sigma033 . Non-standard behavior of $D$ on the damping coefficient $\gamma$ is also possible. A good example is Brownian diffusion of skyrmions within magnetic materials skyrmion1 . For vanishing topological charge, $D$ monotonically decreases with growing $\gamma$. However, if the topological charge is non-zero, the diffusion coefficient increases until it reaches its maximum and next it decreases when the damping constant grows skyrmion2 . In the paper, we study the issue of non-monotonic dependence of diffusion on the friction coefficient for the Brownian particle moving in a spatially periodic potential and driven by a time-periodic force, i.e., the model which is relevant to a variety of different context risken , including pendulums gitterman , Josephson junctions kautz1996 , superionic conductors fulde , dipoles rotating in external fields coffey , Frenkel-Kontorova models frenkel , charge density waves gruner and cold atoms dwelling in optical lattices renzoni ; denisov , to name only a few. We mention that a similar system, however with an additional constant force, has been studied in Ref. march-prl .
The layout of the present work is as follows. In Sec. II, we describe the model of the Brownian particle moving in a spatially periodic potential driven by an unbiased time-periodic force. It is defined in terms of the dimensionless Langevin equation. Its full dimensional version is presented in Appendix A. In Sec. III, we analyze the diffusion coefficient as a function of dissipation, i.e. how $D$ depends on the friction coefficient $\gamma$. We show that there are several regimes in which (i) $D$ monotonically decreases with respect to $\gamma$; (ii) $D$ is non-monotonic and exhibits maximum as a function of $\gamma$; (iii) there are many extrema of $D$ as $\gamma$ grows. In Sec. IV, we implement the vibrational mechanics approximation march-prl which is valid for high frequency of the ac-driving force and examine this approximation for the analysis of diffusion coefficient as a function of damping. Section V provides a summary and conclusions.
II Model of Brownian dynamics
@@
Note that in this scaling the dimensionless mass $m=1$. The parameter $\gamma$ is the friction coefficient and $Q\propto k_{B}T$ is the dimensionless temperature of the system. The coupling of the particle with thermostat is modeled by the $\delta$-correlated Gaussian white noise $\xi(t)$ of zero mean and unit intensity, i.e. $\langle\xi(t)\rangle=0$ and $\langle\xi(t)\xi(s)\rangle=\delta(t-s)$. The starting dimensional equation is presented in Appendix A, where the corresponding scaling and dimensionless parameters are defined.
The complexity of the underlying dynamics with the nonlinear force $f(x)=-\sin{x}$, the time-periodic force $g(t)=a\sin(\omega t+\phi_{0})$ of the amplitude $a$, frequency $\omega$, initial phase $\phi_{0}$ and thermal fluctuations of the intensity $2\gamma Q$ do not allow for a analytical approach which presently is clearly beyond the known mathematical methods. Therefore numerical simulations have been employed spiechowicz2015cpc .
@@
III Diffusion coefficient
Despite its apparent simplicity, the system (3) is extremely versatile, exhibits an impressive diversity of behavior and can successfully describe a number of important and various physical situations. We will be interested only in one aspect of these properties, namely, diffusion and its dependence on dissipation characterized by the friction coefficient $\gamma$. Because the parameters space $\{\gamma,a,\omega,Q\}$ is four dimensional we consider only the selected parameter regime in which the diffusion process in the long time limit is normal and described by a time-independent diffusion coefficient $D$. Hence it can be defined as spiechowicz2016scirep
@@ III Diffusion coefficient
III.1 Diffusion versus friction: monotonic decrease
The exact expression for $D$ is known in several cases simpler than the current one. For the force-free Brownian particle or under the influence of a constant force $F$, i.e. when Eq.(3) assumes the form
$\ddot{x}+\gamma\dot{x}=F+\sqrt{2\gamma Q}\,\xi(t),$ (5)
the diffusion coefficient it given by the Einstein relation,
$D=\frac{Q}{\gamma}.$ (6)
$\ddot{x}+\gamma\dot{x}=-\sin{x}+\sqrt{2\gamma Q}\,\xi(t),$ (7)
$D=\frac{Q}{\gamma}\,B(Q),$ (8)
where $B(Q)$ depends only on the dimensionless temperature $Q$.
$B(Q)=\frac{\sqrt{\pi/Q}}{2I_{0}(1/Q)}\int_{1}^{\infty}\frac{\mbox{e}^{-y/Q}\,dy}{\sqrt{y+1}\,\mathbf{E}(\sqrt{2/(y+1)})},$ (9)
where $I_{0}(x)$ is the modified Bessel function of the first kind and $\mathbf{E}(k)$ is the complete elliptic integral of the second kind specialF .
On the other hand, in the opposite case of overdamped dynamics, i.e. when the inertial term in Eq. (7) vanishes, it assumes the following form lifson ; entropy23
$B(Q)=\frac{1}{I_{0}^{2}(1/Q)}.$ (10)
In the above three cases, the diffusion coefficient $D$ depends inversely on the friction coefficient $\gamma$, i.e., $D\sim 1/\gamma$. In other words, if the damping is stronger then the diffusion coefficient is smaller.
Figure 1: Dimensionless diffusion coefficient $D$ as a function of the amplitude $a$ of the ac-driving for its fixed frequency ${\omega=1.59}$ and two values of the friction coefficient $\gamma$. Temperature reads $Q=0.05$.
@@ III.1 Diffusion versus friction: monotonic decrease
Figure 2: Dependence of the diffusion coefficient $D$ on the friction coefficient $\gamma$ for the ac-driving amplitude $a=6$ and its frequency $\omega=1.59$. The temperature is $Q=0.5$.
@@
When the ac-force is included as in Eq. (3), the analytical form of $D$ is not known and only numerical results are available. In Ref. march-chaos it was shown that for a tailored parameter regime the diffusion coefficient $D$ displays the damped quasi-periodic dependence on the amplitude $a$ of the time-periodic force $g(t)=a\sin{(\omega t+\phi_{0})}$, as an example see Fig. 1. The origin of this phenomenon was explained in Ref. marchPRE23 and is rooted in the complex deterministic structure of the nonlinear dynamics governed by a variety of stable and unstable orbits.
We remind that the running and localized states (periodic orbits) can coexist in this system marchPRE23 . The particle can oscillate in the periodic potential wells and next proceed either forward or backward within one or many temporal periods, or it can permanently move in one direction. There are several types of contribution to the spread of trajectories and in consequence to the diffusion coefficient $D$. The first, which is the largest one, comes from the distance between the running trajectories moving into the opposite directions. The second, also relatively large, is the spread between the running and periodic orbits. The next follows from contributions driven by thermal fluctuation in each of the running and localized states. Clearly, even at high temperature the latter are remarkably smaller than the two previous ones. Therefore it is intuitively expected that the diffusion coefficient $D$ will change considerably especially when the population of running and locked states is significantly modified.
The dependence of the diffusion coefficient $D$ on the friction coefficient $\gamma$ has not been studied in Ref. marchPRE23 . In Fig. 1 we show the influence of $\gamma$ on the behavior of $D$ for low temperature regime $Q=0.05$. It is seen that $D$ decreases as $\gamma$ increases, which portrays the expected standard behavior. So, it is similar to the previous cases presented in Eqs. (6) and (8).
In Fig. 2 we show the characteristics $D$ versus $\gamma$ for a set of fixed parameters. For the driving amplitude $a=6$ and temperature $Q=0.5$ (the first local maximum of $D$ in Fig. 1) one can distinguish three distinct intervals of functional dependence: (i) for damping $\gamma\in(0.01,0.1)$ the diffusion coefficient diminishes like $1/\gamma^{1.27}$; (ii) for $\gamma\in(0.1,8)$, it vanishes like $1/\gamma^{2}$; (iii) for larger $\gamma$ it again decreases like $1/\gamma$. The cases (i) and (ii) show different behavior than this in Eqs. (6) and (8). For the driving amplitude $a=10$ (the first local minimum of $D$ in Fig. 1), one can observe a gentle slope in the vicinity of $\gamma\sim 0.5$ and $1/\gamma$ is detected for $\gamma\gg 1$ (not depicted). The above exemplified monotonic dependence of $D$ on $\gamma$ occurs in extended domain of parameter space, in particular for large values of $\gamma$ and high temperature $Q$. However, this picture is changed radically when temperature $Q$ is significantly lowered.
III.2 Diffusion versus friction: non-monotonic dependence
The case illustrated in Fig. 2 is in fact only a quantitative modification of the monotonic decay $D\sim 1/\gamma$. Much more interesting is the non-monotonic dependence $D(\gamma)$. Such a situation is a qualitatively new phenomenon: the non-monotonicity implies the occurrence of local extrema, i.e. maxima and/or minima. An example of this behavior is visualized in Fig. 3. For the driving amplitude $a=14$ corresponding to the second local maximum of $D$ in Fig. 1 and lower temperature $Q=0.05$, the diffusion coefficient decreases for small values of $\gamma$. Next, it attains the local minimum and starts to increase to the local maximum at $\gamma\approx 0.06$. If $\gamma$ continues to grow, the diffusion coefficient monotonically diminishes to the next local minimum and increases up to the next maximum at $\gamma\approx 3.345$. Further increase of $\gamma$ results in a monotonic decline of $D$ to zero. The main question is why $D$ starts to grow when the friction increases and whether this behavior is exception or can occur in a wider domain of the parameters space. Let us note that the local maxima can be detected for both weak ($\gamma\approx 0.06$) as well as strong damping (${\gamma\approx 3.345}$). However, the mechanism behind this effect is radically different in both cases.
At this place it is worth mentioning that in the system there are localized states $V=0$ (the particle oscillates in one or several potential wells and the motion is bounded in the coordinate space) and running states $V=\pm k\omega,k=1,2,3,...$, where $\omega$ is the dimensionless frequency of the periodic force and the period averaged velocity is defined as
@@ III.2 Diffusion versus friction: non-monotonic dependence
Figure 3: Non-monotonic dependence of the diffusion coefficient $D$ on the friction coefficient $\gamma$ for the ac-driving amplitude $a=14$ (the second maximum in Fig. 1). The frequency of the ac-driving reads $\omega=1.59$ and temperature is $Q=0.05$.
Figure 4: Stationary probability P of the Brownian particle to reside in various states $V=0,V=\pm k\omega\}$ characterizing the deterministic counterpart of the studied setup. The black line corresponds to the sum of probabilities for all running states. The amplitude is $a=14$ and the frequency reads $\omega=1.59$. Temperature is $Q=0.05$.
In the insets of Fig. 3 the probability distribution $P(V)$ for the period-averaged velocity $V$ of the Brownian particle is depicted for the detected local maxima of diffusion coefficient $D$. For the weak damping, in the vicinity of $\gamma=0.06$, the probability distribution has three remarkable maxima, one corresponding to the locked state $V=0$ and two symmetric running states of the averaged velocities $V=\pm\omega$. At non-zero temperature, there are random transitions induced by thermal fluctuations between these states. For one trajectory, the particle can follow the running solution, next for some time interval it resides in the localized state and later it again travels along the running state either in the same or the opposite direction. In this way, for a given trajectory various deterministic solutions are visited in a random sequence. In consequence the deterministic structure of locked and running states impacts the spread of trajectories.
In Fig. 4 we consider the probability P to reside in various states $\{V=0,\pm k\omega\}$ characterizing the deterministic counterpart of the studied setup. It allows us to answer the question about the mechanism standing behind the observed pronounced maximum of the diffusion coefficient $D$ as a function of the friction $\gamma$ for weak damping regime $\gamma\approx 0.06$. The initial decay of this characteristics is clearly related to a decrease of the running states $V=\pm k\omega$ and simultaneous grow of the locked ones $V=0$. Next, there is one-to-one correspondence between the non-monotonic behavior of $D$ and the probability P to reside in the running states. It means that the local growth of the number of running trajectories is responsible for the increase of $D$ despite the fact that the damping $\gamma$ is getting bigger as well.
Figure 5: Probability density $p(x)$ for the position of the particle in the long time regime averaged over the initial conditions $x_{0}$, $v_{0}$ and $\phi_{0}$. The system is deterministic $Q=0$. The potential $U(x)=-\cos{x}$ is depicted schematically for the convenience of the reader. Other parameters are: $a=14$, $\omega=1.59$ and the damping coefficient: $(a)\,\gamma=3.15;$ $(b)\,\gamma=3.21;$ $(c)\,\gamma=3.3;$ $(d)\,\gamma=3.345;$ $(e)\,\gamma=3.36;$ $(f)\,\gamma=3.41$.
As it is well evidenced in Fig. 4 for strong damping regime $\gamma>1$ the running states cease to exist and the period-averaged velocity distribution $P(V)$ exhibits only one maximum corresponding to the locked orbits, c.f. Fig. 3. Yet we can still detect a non-monotonic behavior of the diffusion coefficient $D$ as a function of the friction coefficient $\gamma$ with a well pronounced maximum observed for $\gamma_{M}\approx 3.345$. Therefore the previously outlined mechanism of this effect for low to moderate $\gamma$ is no longer valid in this regime. We shall now explain the origin of this phenomenon which is clearly different than the former one.
@@
We start our consideration assuming the value of $\gamma=3.15<\gamma_{M}$: if $\gamma$ increases then $D$ increases. To explain this effect we first look at the probability density $p(x)$ for the position of the particle in the long time regime. In Fig. 5 we present this quantity for the deterministic trajectories $Q=0$ averaged with respect to the initial position $x_{0}$, velocity $v_{0}$ and phase $\phi_{0}$ distributed uniformly over the intervals $[0,2\pi]$, $[-2,2]$ and $[0,2\pi]$, respectively. We note that the averaging over the phase $\phi_{0}$ in the long time regime is equivalent to the averaging over the period of the external driving. For the convenience of the reader in the same plot we additionally illustrate the periodic potential $U(x)=-\cos{x}$. The reader can observe the characteristic arc-sine form of $p(x)$ levy ; arcsine which here is a consequence of averaging over the phase $\phi_{0}$.
@@
It turns out that in the presented regime corresponding to $\gamma=3.15$ (panel (a) in Fig. 5) the deterministic particle symmetrically oscillates around the maximum $x=x_{M}=\pi$ of the potential and covers the indicated distance $\Delta x$ between its minimal and maximal position. The quantity $\Delta x/2$ can be interpreted as the amplitude of oscillations and the boundary values of this interval correspond to the most probable values of the particle position. Oscillations around the maximum of the potential (the inverted pendulum) are not astonishing taking into account the large amplitude $a=14$ of the time periodic force. Note that the total ”rocking” potential $V(x,t)=-\cos(x)-ax\cos(\omega t+\phi_{0})$ has local extrema (and the barriers) only when $a|\cos(\omega t+\phi_{0})|<1$. It means that for the amplitude $a=14$ in majority of time intervals ($95.5\%$ of full time) there are no barriers. Only in very, very short time itervals ($4.5\%$ of full time) the barriers emerge and the potential is indeed the washboard one. Two extreme forms $V_{1}=-\cos{x}-14x$ and $V_{2}=-\cos{x}+14x$ looks like straight lines, the barriers disappear for some time intervals and the particle can ”dance” around any position.
@@
If now $\gamma$ grows to the value $\gamma_{1}=3.2$, the amplitude of oscillations decreases and the spatial interval $\Delta x$ around the maximum $x_{M}$ of the potential $U(x)$ decreases. It means that in this window the effective potential barrier the particle needs to overcome to diffuse due to thermal agitation is reduced as well. It explains why the diffusion coefficient $D$ increases despite the grow of damping $\gamma$.
@@
Next, at $\gamma_{1}=3.2$ the ”phase transition” arises: two periodic orbits emerge, oscillate around the potential maximum corresponding to the indicated intervals $\Delta x_{1}$ and $\Delta x_{2}$ and they move in the opposite directions when $\gamma$ increases, see panel (c) in Fig. 5. These oscillations are not symmetric around $x_{M}$ but their centers are shifted towards the potential minimum. It means that one of the turning points $x_{e}$ of the orbit (where the particle velocity changes sign) is closer to the potential maximum and the barrier height is further reduced, see the two inner vertical blue and red lines in panel (c) of Fig. 5.
@@
For $\gamma_{M}=3.345$ one of the turning points of both blue and red orbits is exactly at the potential maximum $x_{M}$ and consequently the barrier height is zero, cf. panel (d) in Fig. 5. When the particle is approaching $x_{M}$ its velocity decreases to zero and for a relatively long time it resides in the neighborhood of $x_{M}$. In consequence, thermal fluctuations can freely spread the particle and diffusion is maximal.

### v2

# Giant increase of diffusion by small rise of friction

@@ Abstract
Abstract
Diffusion coefficient usually decreases when friction increases. We analyze the opposite behavior in the paradigmatic system consisting of an inertial Brownian particle moving in a symmetric spatially periodic potential and driven by an unbiased time periodic force. For tailored parameter set in strong dissipation regime the particle spreading can be giantly amplified: if the friction is twice as large then the diffusion grows up to five orders of magnitude. The mechanism lying behind this effect is related to bifurcation of periodic orbits oscillating around the potential maximum and their symmetric displacement towards the adjacent potential minima when the friction coefficient increases. On the other hand, in the weak dissipation regime, where the increase of diffusion vs friction is also observed, the effect is induced by a non-monotonic change of population of the running orbits. However, in this regime the enhancement of diffusion is much smaller.
Introduction
@@
$D=D_{E}=\frac{k_{B}T}{\gamma},$ (1)
where $k_{B}$ is the Boltzmann constant, $T$ is temperature of the medium (thermostat) in which the particle moves and $\gamma$ is the friction (damping) coefficient which characterizes the interaction strength of the Brownian particle with the environment. The same relation is also fulfilled for a Brownian particle subjected to a constant force $F={const}$. From this equation it follows that $D_{E}$ depends inversely on the friction coefficient $\gamma$. In other words, $D_{E}$ decreases when $\gamma$ increases. Studies on e.g. surface diffusion reveals that in some regimes the more general relation is satisfied, namely,
@@
The exponent $\sigma$ is not universal but depends on a system, its model and the corresponding assumptions. Different values have been reported like $\sigma=1/2$ sigma05 , $\sigma=1/3$ sigma03 , $\sigma\in(0.64,0.76)$ sigma064 , $\sigma\in(0,1/3)$ sigma033 . Non-standard behavior of $D$ on the friction coefficient $\gamma$ is also possible. A good example is Brownian diffusion of skyrmions within magnetic materials skyrmion1 . For vanishing topological charge, $D$ monotonically decreases with growing $\gamma$. However, if the topological charge is non-zero, the diffusion coefficient increases until it reaches its maximum and next it decreases when the friction constant grows skyrmion2 . In this paper, we study the issue of non-monotonic dependence of diffusion on the friction coefficient for the Brownian particle moving in a spatially periodic potential and driven by a time-periodic force, i.e., the model which is relevant to a variety of different context risken , including pendulums gitterman , Josephson junctions kautz1996 , superionic conductors fulde , dipoles rotating in external fields coffey , Frenkel-Kontorova models frenkel , charge density waves gruner and cold atoms dwelling in optical lattices renzoni ; denisov , to name only a few. We mention that a similar system, however with an additional constant force, has been studied in Ref. march-prl .
Below we describe in detail the model of the Brownian particle moving in a spatially periodic potential driven by an unbiased time-periodic force and analyze the diffusion coefficient as a function of dissipation, i.e. how $D$ depends on the friction coefficient $\gamma$. We show that there are several distinct regimes in which (i) $D$ monotonically decreases with respect to $\gamma$ or (ii) $D$ is non-monotonic and exhibits a maximum as a function of $\gamma$. Within a tailored parameter set we identify a giant enhancement of the diffusion coefficient when the friction coefficient is slightly increased. We explain the mechanism responsible for this phenomenon. We also apply the analytical but approximate theory of vibrational mechanics and verify its relevance to the studied problem.
Model of Brownian dynamics
@@
Note that in this scaling the dimensionless mass $m=1$. The parameter $\gamma$ is the friction coefficient and $Q\propto k_{B}T$ is the dimensionless temperature of the system. The coupling of the particle with thermostat is modeled by the $\delta$-correlated Gaussian white noise $\xi(t)$ of zero mean and unit intensity, i.e. $\langle\xi(t)\rangle=0$ and $\langle\xi(t)\xi(s)\rangle=\delta(t-s)$. The starting dimensional equation is presented in the section Methods, see Eq. (14), where the corresponding scaling and dimensionless parameters are defined.
The complexity of the underlying dynamics with the nonlinear force $f(x)=-\sin{x}$ and the corresponding spatially periodic potential $W(x)=-\cos(x)$, the time-periodic force $g(t)=a\sin(\omega t+\phi_{0})$ of the amplitude $a$, frequency $\omega$, initial phase $\phi_{0}$ and thermal fluctuations of the intensity $2\gamma Q$ do not allow for an analytical approach which presently is clearly beyond the known mathematical methods. Therefore numerical simulations have been employed spiechowicz2015cpc .
@@
Results: Impact of friction on diffusion
Despite its apparent simplicity, the system (3) is extremely versatile, exhibits an impressive diversity of behavior and can successfully describe a number of important and various physical situations. We will be interested only in one aspect of these properties, namely, diffusion and its dependence on dissipation characterized by the friction coefficient $\gamma$. Because the parameters space $\{\gamma,a,\omega,Q\}$ is four dimensional we consider only the selected parameter regime in which the diffusion process in the long time limit is normal and described by a time-independent diffusion coefficient $D$. Hence it can be defined as jpc
@@
Figure 1: From the weak to strong dissipation regime: The diffusion coefficient $D$ versus the friction coefficient $\gamma$. Parameters are: the ac-driving amplitude $a=6$, its frequency $\omega=1.59$ and temperature $Q=0.5$.
@@
The exact expression for the diffusion coefficient is not known for the model (3) and only numerical results are available. In Ref. march-chaos it was shown that for a tailored parameter regime the diffusion coefficient $D$ displays the damped quasi-periodic dependence on the amplitude $a$ of the time-periodic force $g(t)=a\sin{(\omega t+\phi_{0})}$. The origin of this phenomenon was explained in Ref. marchPRE23 and is rooted in the complex deterministic structure of the nonlinear dynamics governed by a variety of stable and unstable orbits. The dependence of the diffusion coefficient $D$ on the friction coefficient $\gamma$ has not been studied in Ref. marchPRE23 . In Fig. 1 we show expected decrease of $D$ versus $\gamma$ for a set of some fixed parameters. For the driving amplitude $a=6$ and temperature $Q=0.5$ one can distinguish three distinct intervals of functional dependence: (i) for friction $\gamma\in(0.01,0.1)$ the diffusion coefficient diminishes like $1/\gamma^{1.27}$; (ii) for $\gamma\in(0.1,8)$, it decays like $1/\gamma^{2}$; (iii) for larger $\gamma$ it again decreases like $1/\gamma$. The above exemplified monotonic dependence of $D$ on $\gamma$ occurs in extended domain of the parameter space. However, this picture is changed radically when temperature $Q$ is significantly lowered. We have searched a part of the three dimensional parameter space $\{\gamma,a,Q\}$ in order to locate regions of non-monotonic dependence of $D$ on $\gamma$.
We remind that in the noiseless system $Q=0$ corresponding to Eq. (3), the dynamics may become complex, exhibiting both regular and chaotic behavior kissner in the four-dimensional parameter space $\{\gamma,a,\omega,Q\}$ in which running and localized states (periodic orbits) can coexist marchPRE23 . The particle can oscillate in the periodic potential wells and next proceed either forward or backward within one or many temporal periods, or it can permanently move in one direction. Depending upon the system parameters and initial conditions various stationary regimes can be described by a rough approximation
$x(t)=x_{c}+G(\omega t)\pm k\omega t,$ (5)
where $G(\omega t)$ is a periodic function of time and characterizes oscillations, the parameter $x_{c}$ is a center of oscillations and $\omega$ is the dimensionless frequency of the time periodic force. In the case of locked trajectories $k=0$ and for running solutions $k\neq 0$. For the mechanical pendulum, $k=0$ corresponds to oscillations around $x_{c}$. The case $k\neq 0$ corresponds to rotation of the pendulum. In our recent work marchPRE23 we exemplified the system parameter sets for which locked and running solutions can emerge. Generally, it is difficult to establish a relationship between values of the parameters and initial conditions to answer the question for which of their subsets running or locked trajectories can be observed. However, our intuition suggests that for a given parameter set and for sufficiently large $\gamma$ there are only locked states.
Due to symmetry of the system (3), in the long time limit the directed velocity $\langle v\rangle$ must vanish for both zero and non-zero temperature regimes,
$\langle v\rangle=0.$ (6)
@@
For the noisy system with $Q>0$, there are several types of contribution to the spread of trajectories and in consequence to the diffusion coefficient $D$. The first, which is the largest one, comes from the distance between the running trajectories moving into the opposite directions. The second, also relatively large, is the spread between the running and periodic orbits. The next follows from contributions driven by thermal fluctuation in each of the running and localized states. Clearly, even at high temperature the latter are remarkably smaller than the two previous ones. Therefore it is intuitively expected that the diffusion coefficient $D$ will change considerably especially when the population of running and locked states is significantly modified. An example of various types of trajectories is visualized in Fig. 4 in Ref. marchPRE23 .
A. Giant enhancement of diffusion
The case illustrated in Fig. 1 is in fact only a quantitative modification of the monotonic decay $D\sim 1/\gamma$. Much more interesting is the non-monotonic dependence $D(\gamma)$. Such a situation is a qualitatively new phenomenon: the non-monotonicity implies the occurrence of local extrema, i.e. maxima and/or minima. An example of this behavior is visualized in Fig. 2. For the driving amplitude $a=14$, its frequency $\omega=1.59$ and lower temperature $Q=0.05$, the diffusion coefficient decreases for small values of $\gamma$ and attains the local minimum $D=2.5\times 10^{-7}$ at $\gamma=1.66$. Next, it starts to increase to the local maximum $D=1.4\times 10^{-2}$ at $\gamma=\gamma_{M}\approx 3.3$. Further increase of $\gamma$ results in a monotonic decline of $D$ to zero. We note that the enhancement of $D$ is $5.6\times 10^{4}$, i.e. almost by five orders of magnitude when friction is increased less than twice. The main question is why $D$ starts to increase when the friction grows and whether this behavior is exception or can occur in a wider domain of the parameters space. We shall now explain the origin of this phenomenon. In the deterministic counterpart of the noisy system there are localized states with $V=0$ and running states with $V=\pm k\omega,k=1,2,3,...$. The period averaged velocity is defined as
[…]
@@ A. Giant enhancement of diffusion
Figure 2: The strong dissipation regime: Non-monotonic dependence of the diffusion coefficient $D$ on the friction coefficient $\gamma$. Inset: The probability distribution $P(V)$ for the period-averaged velocity $V$ of the Brownian particle. Parameters are: the ac-driving amplitude $a=14$, its frequency $\omega=1.59$ and temperature $Q=0.05$.
Figure 3: The probability density $p(x)$ for position of the particle in the long time regime averaged over the initial conditions $x_{0}$, $v_{0}$ and $\phi_{0}$. The system is deterministic $Q=0$. Black curves: the potential $W(x)=-\cos{x}$. Other parameters are: $a=14$, $\omega=1.59$ and the friction coefficient: $(a)\,\gamma=3.15;$ $(b)\,\gamma=3.21;$ $(c)\,\gamma=3.3;$ $(d)\,\gamma=3.345;$ $(e)\,\gamma=3.36;$ $(f)\,\gamma=3.41$. For $\gamma>3.2$, in dependence of initial conditions, two classes of periodic orbits emerge which are symmetrically shifted in the positive (blue) and negative (red) directions from the value $x_{M}=\pi$. Bottom: The pitchfork bifurcation: if $\gamma<3.2$ the centers $x_{c}$ of periodic orbits are located at the maximum $x=\pi$ of the potential $W(x)$ and for $\gamma>3.2$ they are shifted to its adjacent minima $x=0$ and $x=2\pi$.
@@
In the insets of Fig. 2 the probability distribution $P(V)$ for the period-averaged velocity $V$ of the Brownian particle is depicted for the detected maximum of the diffusion coefficient $D$. In the regime of strong damping $\gamma\sim 3$ the running states are short lived and the period-averaged velocity distribution $P(V)$ exhibits only one maximum corresponding to the locked orbits ($k=0$ in Eq. (5)). This fact in turn is important for explanation of the giant enhancement of $D$. We first look at the probability density $p(x)$ for the position of the particle in the long time regime. In Fig. 3 we present this quantity for the deterministic trajectories $Q=0$ averaged with respect to the initial position $x_{0}$, velocity $v_{0}$ and phase $\phi_{0}$ distributed uniformly over the intervals $[0,2\pi]$, $[-2,2]$ and $[0,2\pi]$, respectively. We note that the averaging over the phase $\phi_{0}$ in the long time regime is equivalent to the averaging over the period of the external driving. For the convenience of the reader in the same plot we additionally illustrate the periodic potential $W(x)=-\cos{x}$. The reader can observe the characteristic arc-sine form of $p(x)$ levy ; arcsine which here is a consequence of averaging over the phase $\phi_{0}$.
@@
We start our consideration assuming the value of $\gamma=3.15<\gamma_{M}$: if $\gamma$ increases then $D$ increases. It turns out that in the presented regime (panel (a) in Fig. 3) the deterministic particle symmetrically oscillates around the maximum $x=x_{M}=\pi$ of the potential and covers the indicated distance $\Delta x$ between its minimal and maximal position. The quantity $\Delta x/2$ can be interpreted as the amplitude of oscillations and the boundary values of this interval correspond to the most probable values of the particle position. In this case the center of oscillations is $x_{c}=\pi$ and oscillations around the maximum of the potential (the inverted pendulum) are not astonishing taking into account the large amplitude $a=14$ of the time periodic force. Note that the total ”rocking” potential $\mathbb{W}(x,t)=-\cos(x)-ax\cos(\omega t+\phi_{0})$ has local extrema (and the barriers) only when $a|\cos(\omega t+\phi_{0})|<1$. Therefore for the amplitude $a=14$ in majority of time intervals ($95.5\%$ of full time) there are no barriers. Only in short time intervals ($4.5\%$ of full time) the barriers emerge and the potential is indeed the washboard one. Two extreme forms $\mathbb{W}_{1}=-\cos{x}-14x$ and $\mathbb{W}_{2}=-\cos{x}+14x$ looks like straight lines, the barriers disappear for some time intervals and the particle can ”dance” around any position.
@@
If now $\gamma$ grows to the value $\gamma_{1}=3.2$, the amplitude of oscillations decreases and the spatial interval $\Delta x$ around the maximum $x_{M}$ of the potential $\mathbb{W}(x)$ decreases. In consequence, in this window the effective potential barrier, the particle needs to overcome to diffuse due to thermal agitation, is reduced as well. It explains why the diffusion coefficient $D$ can increase despite the growth of friction $\gamma$. We should also remember that in the noisy system the intensity $\gamma Q$ of thermal noise $\xi(t)$ also increases as $\gamma$ increases and it is the second factor which can lead to the increase of $D$.
@@
Next, at $\gamma_{1}=3.2$ the ”phase transition” arises: In dependence of initial conditions, two classes of periodic orbits emerge with two values of $x_{c}=x_{c1}$ and $x_{c}=x_{c2}$ which are symmetrically shifted in the positive and negative directions from the value $x_{M}=\pi$. When $\gamma$ increases $x_{c1}$ and $x_{c2}$ move towards $x=0$ and $x=2\pi$, respectively, see panels (b) and (c) in Fig. 3. These oscillations are not symmetric around $x_{M}$ but their centers are shifted towards the potential minima. It means that one of the turning points $x_{e}$ of the orbit (where the particle velocity changes sign) is closer to the potential maximum and the barrier height is further reduced, see the two inner vertical blue and red lines in panel (c) of Fig. 3.
@@
For $\gamma_{M}=3.345$ one of the turning points of both blue and red orbits is exactly at the potential maximum $x_{M}$ and consequently the barrier height is zero, cf. panel (d) in Fig. 3. When the particle is approaching $x_{M}$ its velocity decreases to zero and for a relatively long time it resides in the neighborhood of $x_{M}$. In consequence, thermal fluctuations can freely spread the particle.